OpenStudy (anonymous):

Scott earns $5 for each lawn he mows. On Saturday he mowed 6 lawns, and on Sunday he mowed 8 lawns. Which of the following expressions represents how much Scott earned for the weekend (in dollars)?

OpenStudy (anonymous):

6 * 5 = 30. (For Saturday) 6 * 8 = 48 (For Sunday)

OpenStudy (anonymous):

6(5)+8(5)=$$ for the weekend